OpenStudy (ebydaqueen19):

Help Help!Find the unit rate 4/6=_/1

jimthompson5910 (jim_thompson5910):

in the denominator, they divided by 6 to get that 1 so you have to do the same in the numerator 4/6 = 2/3 = 0.667 (roughly)

jimthompson5910 (jim_thompson5910):

so that's why 4/6 = 0.667/1 (approximately)

OpenStudy (ebydaqueen19):

the problem is about mowing lawns 4 lawns in 6 hours anfd they also want to know how long it will take the person to mow 15 lawns

jimthompson5910 (jim_thompson5910):

4 lawns/6 hours = 15 lawns/x hours 4/6 = 15/x 4x = 6*15 4x = 90 x = ???

OpenStudy (ebydaqueen19):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (unknownrandom):

Another easy way to do this is to cross multiply. Set up 4/6=15/? So 6*15=90 90/4=22.5

OpenStudy (ebydaqueen19):

thanks @UnknownRandom

OpenStudy (ebydaqueen19):

im sorry i still dont know how to find the unit rate though

OpenStudy (unknownrandom):

No problem @ebydaqueen19

OpenStudy (ebydaqueen19):

they said the problem is 4 lawns/hours / _lawns/1 hour

OpenStudy (ebydaqueen19):

never mind